OpenStudy (firejay5):

Solve Logarithm equation. Show work and explain. 30. log5 (x^2 + 7) = 2/3log5 64

OpenStudy (firejay5):

@Hero

Nnesha (nnesha):

same like previous question try it first

OpenStudy (firejay5):

I did try it out and got x = -3 which is wrong

Nnesha (nnesha):

show your work so i canlook over

OpenStudy (firejay5):

haha no trust me plug -3 into the equation you get different answers

Nnesha (nnesha):

it doesn't have to do anything about trust i just said it bec i want to see where you did a mistake so we don't have to start it again that's it

OpenStudy (firejay5):

wait hold on

OpenStudy (firejay5):

it's right I had it right? (medal)

Nnesha (nnesha):

right what ?? -3 ?? how did you get that ?

OpenStudy (firejay5):

Really............... x^2 + 7 = 64^2/3 sqrt of x^2 + 7 = sqrt 16 x+ 7 = 4 x = -3

Nnesha (nnesha):

 why did you take square of x^+7 and at the end where did the square root of 7 gone ?